4. Determine the voltages across and current each resistor shown in the circuit below. 1. E= 18 V, r = 0 60 b

Help solve
Answer #1

between point c and b, the two resistance are in parallel. hence their net resistance is given by

1/R = 1/6+1/3

hence R = 2 ohm

These resistances are connected in series with 4-ohm resistance.

hence total resistance of the circuit = 2+4 = 6 ohm

hence total current flowing through the circuit = I = V/R = 18/6 = 3 Amp.

Now, current flowing through 4 ohm resistance = 3 Amp

and the voltage drop across the 4-ohm resistance = V1 = 3*4 = 12 Volt

so voltage drop across c to b = 18-12= 6 Volt

Since 6-ohm and 3-ohm resistnace are in parallel. hence they have same voltage drop.

so both 6-ohm and 3-ohm resistance have voltage = 6 volt

So, Now current in 6-ohm resistance = 6 volt/6ohm = 1 Amp

current in 2-ohm resistance = 6 volt/3ohm = 2Amp